# A JWST Project on 47 Tucanae: Kinematics, energy equipartition and anisotropy of multiple populations

### v1

@@ 3 Stellar Masses
To analyze the mass dependence of the velocity dispersion, we estimated stellar masses for our various datasets using isochrone fitting. We employed isochrones from the Dartmouth Stellar Evolution Database (Dotter et al. 2008) with the following parameters: $\mathrm{[Fe/H]}=-0.75$ dex, age $=$ 13 Gyr, helium mass fraction Y = 0.246, $\alpha$ enhancement [$\alpha$/Fe] $=+0.4$ dex, foreground reddening $E(B-V)=0.03$ mag, and distance modulus $(m-M)_{0}=13.38$ mag, as in Milone et al. (2023). As an example, Figure 6 shows the CMD for Field A, constructed from HST photometry. Stars are color-coded by their estimated mass, and the fitted isochrone is overlaid in black.
Figure 6: $m_{F814W}$ vs. $m_{F606W}-m_{F814W}$ CMD of stars in Field A, color-coded by mass. The black line represents the isochrone used to fit the cluster (see text for details).
@@
HB stars in 47 Tucanae have masses of approximately 0.65 $M_{\odot}$ (Tailo et al. 2020). Given that stars do not undergo significant mass loss during their HB evolution, AGB stars are assumed to have approximately the same mass. However, the mass loss that RGB stars experience before reaching the HB occurs on timescales much shorter than the two-body relaxation time. Consequently, these stars are expected to retain kinematic properties characteristic of higher-mass stars. Thus, we assigned these stars the same mass as the maximum mass observed in our RGB sample (as in Watkins et al. 2022).
4 Anisotropy
@@
To analyze the internal dynamics of multiple populations in 47 Tucanae, we combined our proper motion dataset (Section 2.1) with the population classification described in Section 2. Following the methodology of Cordoni et al. (2024), we converted stellar coordinates and proper motion components ($\alpha$, $\delta$, $\mu_{\alpha}\cos\delta$, $\mu_{\delta}$) to Cartesian coordinates (x, y, $\mu_{x}$, $\mu_{y}$) using Equation 2 from Gaia Collaboration et al. (2020). To analyze stellar motions relative to the cluster for the Gaia dataset, we subtracted the cluster’s motion as determined by Vasiliev & Baumgardt (2021). Subsequently, we projected the proper motions into radial ($\mu_{\mathrm{R}}$) and tangential ($\mu_{\mathrm{T}}$) components with respect to the cluster center. We performed a correction in absolute $\mu_{\mathrm{R}}$ proper motions to account for perspective contraction/expansion due to the bulk motion of 47 Tucanae along the line of sight (Equation 4 of Bianchini et al. 2018).
@@
In the top panel of Figure 7, we compared the mean radial profiles of $\mu_{\mathrm{T}}$ for 1G (blue points) and 2G (purple squares) stars, given that $\mu_{\mathrm{T}}$ serves as a tracer of rotation. For this analysis, we considered Gaia proper motions matched with the photometric catalog of Lee (2022), which covers a wide radial range, since HST and JWST proper motions are relative and do not display signatures of rotation. We binned the data considering equal-frequency bins, with a total of 1712 2G stars and 868 1G stars. Any discrepancies between these profiles could potentially indicate distinct rotational characteristics between the two populations. We did not observe significant differences between the profiles of 1G and 2G stars (see Section 4.2 for a discussion on the velocity dispersion profiles). The bottom panel shows the radial profile of $\mu_{\mathrm{T}}(\sigma_{\mathrm{R}}^{2}+\sigma_{\mathrm{T}}^{2})^{-1/2}$ for 1G and 2G stars, quantifying the ratio of rotational support to random motion pressure. From this analysis, differences in the rotational behavior of 1G and 2G stars become more pronounced, especially in the cluster’s outer regions. Recent studies (Leitinger et al. 2024; Dalessandro et al. 2024) have further explored the rotational kinematics of multiple populations, revealing that 2G stars tend to rotate more rapidly than 1G in 47 Tucanae (see also Bellini et al. 2017, for a general study on the rotation of 47 Tucanae).
We investigated the skewness of the $\mu_{T}$ versus $\mu_{R}$ distribution for 1G and 2G stars. The top right panel of Figure 5 clearly illustrates that the distribution exhibits skewness in the $\mu_{T}$ values. Indeed, a skew in 47 Tucanae was previously reported by Heyl et al. (2017), who analyzed skewness across all stars. In this study, we focused on the skewness of 1G and 2G stars separately to determine whether their orbital properties differ.
Figure 8 presents the skewness in $\mu_{T}$ for 1G stars (depicted as blue points) and 2G stars (shown as purple points) in the top panel. To compute the skewness, we divided the data into equal star count bins based on radial distance. For each bin, we used bootstrap resampling to estimate the skewness. Specifically, we generated random samples from the data by resampling with replacement and assuming normal distributions based on the observed data and their associated errors. We then computed the skewness for each bootstrap sample. The mean skewness value for each bin is shown, with the standard deviation indicating the uncertainty. Open symbols refer to Gaia data, while filled symbols represent HST/JWST data. Dashed (Gaia) and solid (HST/JWST) lines correspond to the mean skewness for each population determined by calculating the skewness using the proper motions of all stars of each population at all radii. We display 1G measurements in blue, and 2G in purple. The bottom panel displays analogous results for $\mu_{R}$. On average, both 1G and 2G stars exhibit approximately zero skewness in $\mu_{R}$. However, in the tangential component, we observed an average skewness significantly higher for 1G stars compared to 2G stars. These results are consistent with the measurements of Heyl et al. (2017) for skewness of all stars combined in 47 Tucanae. These findings suggest distinct orbital behaviors between 1G and 2G stars within the cluster.
@@
Prior to analyzing the velocity dispersion, anisotropy profiles, and energy equipartition, we removed the rotational component from the Gaia proper motions. This was achieved by fitting sinusoidal functions to the proper motions in the $x$ and $y$ directions as a function of stellar position angle. The residuals from these fits were then adopted as the de-rotated proper motions in the respective directions.
To determine the velocity dispersion, we divided the field of view into a series of concentric annular regions, each including approximately the same number of stars. For each annulus we determined the dispersion from either Gaia DR3, JWST or HST radial and tangential velocity components by maximizing the log-likelihood function described in Equation 1 of Libralato et al. (2022), for HST/JWST, and Bianchini et al. (2018)111updated including the covariance term as in Sollima et al. (2019) for Gaia. We determined the uncertainties in the velocity dispersions using the Markov Chain Monte Carlo algorithm emcee (Foreman-Mackey et al. 2013). Figure 9 illustrates the velocity dispersion and anisotropy (as defined in Equation 1) profiles for all stars in our dataset. Our observations include proper motions of white dwarf stars in Field A, that are highlighted in turquoise.
@@ 4.1 Velocity anisotropy model
Following the results of simulations of the evolution of multiple-population GCs, we model the 2D radial variation of the velocity anisotropy using the following analytical expression (see Dalessandro et al. 2024, and Aros et al. 2025, submitted).
@@
where $\sigma_{\mathrm{T}}$ and $\sigma_{\mathrm{R}}$ are the velocity dispersions in the tangential and radial directions, respectively, $R_{\mathrm{a}}$ is the anisotropy radius, $\beta_{\infty}$ is the anisotropy at $R\gg R_{\mathrm{a}}$, and $R_{\mathrm{t}}$ is the truncation radius, defined as the radius for which the velocity becomes isotropic again (the model assumes isotropy at the cluster center, and increasing anisotropy for $R>R_{\mathrm{a}}$; see Dalessandro et al. 2024, and Aros et al. 2025, submitted, for more details). This parametrization is a modification of the Osipkov-Merrit model (Osipkov 1979; Merritt 1985), allowing a free parameter for the anisotropy (either radial or tangential velocity dispersion for $R>R_{\mathrm{a}}$).
@@
4.2 Kinematic profiles
Figure 10 shows the velocity dispersion profiles in the radial (top panels) and tangential (middle panels) directions, along with the anisotropy profiles (bottom panels). The best-fit model profiles are indicated by the dark solid lines. The first column shows 1G stars (in blue), while 2G stars are shown in purple in the second column. For better comparison, the third column displays the comparison between the 1G and 2G best fit profiles, obtained as described in section 4.1. The radial distances are expressed in units of half-light radius ($R_{\mathrm{h}}$), which is 3.17 arcminutes for 47 Tucanae (Harris 1996, 2010 version).
1G and 2G stars present similar velocity dispersions in the radial direction. In the tangential direction, 1G stars present higher velocity dispersion than 2G stars, except in the very center and outer regions (approximately $\geq 5$ $R_{h}$), where they are very similar. This difference translates into different anisotropy profiles. While 1G stars are consistent with being isotropic, 2G stars present radial anisotropy. Overall, Gaia and HST/JWST profiles are in good agreement within the uncertainties. However, since Gaia targets giant stars, while HST and JWST proper motions are for MS stars, small differences are expected due to energy equipartition (see e.g. Pavlík et al. 2024).
@@ 5 Energy Equipartition
$\small\begin{split}L(\mu_{R},\mu_{T},m\mid\Theta)=&\prod_{i}^{N}(2\pi(\sigma(m_{i}\mid\Theta)^{2}+\delta_{\mathrm{\mu_{R}},i}^{2}))^{-1/2}(2\pi(\sigma(m_{i}\mid\Theta)^{2}+\delta_{\mathrm{\mu_{T}},i}^{2}))^{-1/2}\\ &\times\exp\Bigg{(}-\frac{1}{2}\frac{{\mathrm{\mu_{R}},i}^{2}}{\sigma(m_{i}\mid\Theta)^{2}+\delta_{\mathrm{\mu_{R}},i}^{2}}-\frac{1}{2}\frac{{\mathrm{\mu_{T}},i}^{2}}{\sigma(m_{i}\mid\Theta)^{2}+\delta_{\mathrm{\mu_{T}},i}^{2}}\Bigg{)},\end{split}$ (9)
@@
Figure 12 displays the total velocity dispersion as a function of stellar mass for the fields illustrated in Figure 1. The top row shows the fits of the degrees of energy equipartition $\eta$ and $\mu$ for 1G stars, while the second row shows the same for 2G stars. The uncertainties are derived from the 16th and 84th percentiles of the sampled parameter distributions, representing the lower and upper bounds of a 68% credible interval. We present the results separately for each field, as they have different radial coverage, and energy equipartition can vary with radius, exhibiting higher values in the center of the cluster (Bianchini et al. 2018). The average radius for each field is indicated within the subplots. Each bin contains more than 100 stars, with denser regions containing approximately 400 stars per bin. We note that the bins are displayed for visual comparison only and not for fitting purposes.
@@ 6 Summary and discussion
  • We compared the radial profiles of mean $\mu_{T}$ for 1G and 2G stars by matching Gaia DR3 with the photometric catalog of Lee (2022), which covers an area of $1^{\circ}\times 1^{\circ}$. We did not find significant differences in these profiles. However, our analysis shows that 2G stars have larger values of the ratio of rotational velocity to velocity dispersion indicating differences in the relative strength of rotational to the random motion components. Moreover, we measured skewness for the proper motion components, finding that 1G stars exhibit higher skewness in $\mu_{\mathrm{T}}$ compared to 2G stars, further indicating that these populations differ in their orbital properties.
@@
  • Deep multi-epoch HST observations allowed us to determine proper motions of white dwarf stars in field A. Since multiple populations could not be identified among white dwarfs, we compared the kinematics of these stars with the overall dynamical profiles of all stars in other evolutionary stages in the cluster. We found no significant differences between the kinematics of white dwarfs and those of the remaining stars.

### v2

@@ 4 Stellar masses
To analyze the mass dependence of the velocity dispersion, we estimated stellar masses for our various datasets using isochrone fitting. We employed isochrones from the Dartmouth Stellar Evolution Database (Dotter et al. 2008) with the following parameters: $\mathrm{[Fe/H]}=-0.75$ dex, age $=$ 13 Gyr, helium mass fraction Y = 0.246, $\alpha$ enhancement [$\alpha$/Fe] $=+0.4$ dex, foreground reddening $E(B-V)=0.03$ mag, and distance modulus $(m-M)_{0}=13.38$ mag, as in Milone et al. (2023b). As an example, Figure 6 shows the CMD for Field A, constructed from HST photometry. Stars are color-coded by their estimated mass, and the fitted isochrone is overlaid in black.
Figure 6: $m_{\rm F814W}$ vs. $m_{\rm F606W}-m_{\rm F814W}$ CMD of stars in Field A. The pink line represents the isochrone used to fit the cluster (see text for details), and the horizontal dotted lines indicate initial MS mass values along the CMD inferred from the isochrone.
@@
HB stars in 47 Tucanae have masses of approximately 0.65 $M_{\odot}$ (Tailo et al. 2020). Given that stars do not undergo significant mass loss during their HB evolution, AGB stars are assumed to have approximately the same mass. However, the mass loss that RGB stars experience before reaching the HB occurs on timescales much shorter than the two-body relaxation time. Consequently, AGB and HB stars are expected to retain kinematic properties characteristic of higher-mass stars. Thus, we assigned these stars the maximum mass measured in our RGB sample (as in Watkins et al. 2022).
5 Internal kinematics
@@
To analyze the internal dynamics of the multiple populations in 47 Tucanae, we combined our proper motion dataset (Section 3) with the population classification described in Section 2. Following the methodology of Cordoni et al. (2025), we converted stellar coordinates and proper motion components ($\alpha$, $\delta$, $\mu_{\alpha}\cos\delta$, $\mu_{\delta}$) to Cartesian coordinates (x, y, $\mu_{x}$, $\mu_{y}$) using Equation 2 from Gaia Collaboration et al. (2018). To analyze stellar motions relative to the cluster for the Gaia dataset, we subtracted the cluster’s motion as determined by Vasiliev & Baumgardt (2021). Subsequently, we projected the proper motions into radial ($\mu_{\mathrm{R}}$) and tangential ($\mu_{\mathrm{T}}$) components with respect to the cluster center. We performed a correction in absolute $\mu_{\mathrm{R}}$ proper motions to account for perspective contraction/expansion due to the bulk motion of 47 Tucanae along the line of sight (Equation 4 of Bianchini et al. 2018a).
5.1 Rotation
@@
In the top panel of Figure 7, we compared the mean radial profiles of $\mu_{\mathrm{T}}$ for 1G (blue points) and 2G (purple squares) stars, given that $\mu_{\mathrm{T}}$ serves as a tracer of rotation. For this analysis, we considered Gaia proper motions matched with the photometric catalog of Lee (2022), which covers a wide radial range, since HST and JWST proper motions are relative and do not display signatures of rotation. We binned the data considering equal-frequency bins, with a total of 1712 2G stars and 868 1G stars. Any discrepancies between these profiles could potentially indicate distinct rotational characteristics between the two populations. We did not observe significant differences between the profiles of 1G and 2G stars (see Section 5.5 for a discussion on the velocity dispersion profiles). The bottom panel shows the radial profile of $\mu_{\mathrm{T}}(\sigma_{\mathrm{R}}^{2}+\sigma_{\mathrm{T}}^{2})^{-1/2}$ for 1G and 2G stars, quantifying the ratio of rotational support to random motion pressure. From this analysis, differences in the rotational behavior of 1G and 2G stars become more pronounced, especially in the cluster’s outer regions. Recent studies (Leitinger et al. 2025; Dalessandro et al. 2024) have further explored the rotational kinematics of multiple populations, revealing that 2G stars tend to rotate more rapidly than 1G in 47 Tucanae (see also Bellini et al. 2017, for a general study on the rotation of 47 Tucanae).
@@
5.2 Skewness
We investigated the skewness of the $\mu_{T}$ versus $\mu_{R}$ distribution for 1G and 2G stars. Skewness quantifies the asymmetry of a probability distribution about its mean. We measured skewness using the Fisher-Pearson coefficient (implemented via the skew function in SciPy), which is the third standardized moment of the distribution. A positive skewness indicates a distribution with an elongated tail toward values larger than the mean, while negative skewness indicates a tail extending toward values smaller than the mean. The top right panel of Figure 5 clearly illustrates that the distribution exhibits skewness in the $\mu_{T}$ values. Indeed, a skew in 47 Tucanae was previously reported by Heyl et al. (2017), who analyzed skewness across all stars. In this study, we focused on the skewness of 1G and 2G stars separately to determine whether their orbital properties differ.
Figure 8 presents the skewness in $\mu_{T}$ for 1G stars (depicted as blue points) and 2G stars (shown as purple points) in the top panel. To compute the skewness, we divided the data into equal star count bins based on radial distance. For each bin, we used bootstrap resampling to estimate the skewness. Specifically, we generated random samples from the data by resampling with replacement and assuming normal distributions based on the observed data and their associated errors. We then computed the skewness for each bootstrap sample. The mean skewness value for each bin is shown, with the standard deviation indicating the uncertainty. Open symbols refer to Gaia data, while filled symbols represent HST/JWST data. Dashed (Gaia) and solid (HST/JWST) lines correspond to the mean skewness for each population determined by calculating the skewness using the proper motions of all stars of each population at all radii. We display 1G measurements in blue, and 2G in purple. The bottom panel displays analogous results for $\mu_{R}$. On average, both 1G and 2G stars exhibit approximately zero skewness in $\mu_{R}$. However, in the tangential component, we observed an average skewness significantly higher for 1G stars compared to 2G stars. These results are consistent with the measurements of Heyl et al. (2017) for skewness of all stars combined in 47 Tucanae.
@@ 5.2 Skewness
In studies of galaxy kinematics based on line-of-sight velocities, skewed velocity distribution profiles are often linked to internal rotation (see, e.g., van der Marel & Franx 1993; van de Sande et al. 2017). However, in our investigation — based on proper motion measurements — the interpretation of differences (or lack thereof) in skewness between the tangential and radial velocity distributions is less straightforward. Notably, the differences in tangential velocity skewness between the two stellar generations do not correspond to significant differences in their rotational velocities.
Other dynamical effects could also play a role in shaping the velocity distribution profile. In particular, the complex interaction between internal dynamical processes and the Galactic tidal field along with the escape of stars could affect the skewness of the cluster’s velocity distribution. Any effect related to the external tidal field, however, would presumably affect more significantly the cluster’s outermost regions and lead to a radial variation in skewness which contrasts with the observed approximately constant skewness for 1G stars. Therefore, our results do not provide strong evidence that tidal stripping plays a significant role in explaining the observed skewness of 1G and 2G stars. Further exploration through dedicated N-body simulations is needed to better understand how the interplay between the initial dynamical properties of the two populations and the effects of various dynamical processes influences the evolution of the velocity distribution skewness. Nevertheless, the velocity distribution skewness measured in our analysis provides additional evidence of kinematic differences between the two stellar generations and sets a new constraint for models of the formation and dynamical evolution of multiple populations.
5.3 Velocity dispersion
Prior to analyzing the velocity dispersion, anisotropy profiles, and energy equipartition, we removed the rotational component from the Gaia proper motions for consistency with the relative proper motions from HST and JWST. This was achieved by fitting sinusoidal functions to the proper motions in the $x$ and $y$ directions as a function of stellar position angle. The residuals from these fits were then adopted as the de-rotated proper motions in the respective directions.
To determine the velocity dispersion, we divided the field of view into a series of concentric annular regions, each including approximately the same number of stars. For each annulus we determined the dispersion from either Gaia DR3, JWST or HST radial and tangential velocity components by maximizing the log-likelihood function described in Equation 1 of Libralato et al. (2022), for HST/JWST, and Bianchini et al. (2018a)444updated including the covariance term as in Sollima et al. (2019) for Gaia. We determined the uncertainties in the velocity dispersions using the Markov Chain Monte Carlo algorithm emcee (Foreman-Mackey et al. 2013). Figure 9 illustrates the velocity dispersion and anisotropy (as defined in Equation 1) profiles for all stars in our dataset. Our observations include proper motions of white dwarf stars in Field A, that are highlighted in turquoise.
@@ 5.4 Velocity anisotropy model
Following the results of simulations of the evolution of multiple-population GCs, we model the 2D radial variation of the velocity anisotropy using the following analytical expression (see Dalessandro et al. 2024; Aros et al. 2025).
@@
where $\sigma_{\mathrm{T}}$ and $\sigma_{\mathrm{R}}$ are the velocity dispersions in the tangential and radial directions, respectively, $R_{\mathrm{a}}$ is the anisotropy radius, $\beta_{\infty}$ is the anisotropy at $R\gg R_{\mathrm{a}}$, and $R_{\mathrm{t}}$ is the truncation radius, defined as the radius for which the velocity becomes isotropic again (the model assumes isotropy at the cluster center, and increasing anisotropy for $R>R_{\mathrm{a}}$; see Dalessandro et al. 2024; Aros et al. 2025, for more details). This parametrization is a modification of the Osipkov-Merrit model (Osipkov 1979; Merritt 1985), allowing a free parameter for the anisotropy (either radial or tangential velocity dispersion for $R>R_{\mathrm{a}}$).
@@
5.5 Anisotropy of the multiple populations
Figure 10 shows the velocity dispersion profiles in the radial (top panels) and tangential (middle panels) directions, along with the anisotropy profiles (bottom panels). The best-fit model profiles are indicated by the dark solid lines. Triangles correspond to data from fields A (lower-MS), B and C, circles to field D, squares to field A (upper-MS), and crosses to the inner field, JWL, Stetson, and Gaia XP data. The first column shows 1G stars (in blue), while 2G stars are shown in purple in the second column. For better comparison, the third column displays the comparison between the 1G and 2G best fit profiles, obtained as described in section 5.4. The radial distances are expressed in units of half-light radius ($R_{\mathrm{h}}$), which is 3.17 arcminutes for 47 Tucanae (Harris 1996, 2010 version).
1G and 2G stars present similar velocity dispersions in the radial direction. In the tangential direction, 1G stars present higher velocity dispersion than 2G stars, except in the very center and outer regions (approximately $\geq 5$ $R_{h}$), where they are very similar. This difference translates into different anisotropy profiles. While 1G stars’ motions are isotropic, 2G stars present radial anisotropy. Overall, Gaia and HST/JWST profiles are in good agreement within the uncertainties. However, since Gaia targets giant stars, while HST and JWST proper motions are for MS stars, small differences are expected due to energy equipartition (see e.g. Pavlík et al. 2024).
@@ 5.6 Energy equipartition
$\small\begin{split}L(\mu_{R},\mu_{T},m\mid\Theta)=&\prod_{i}^{N}(2\pi(\sigma(m_{i}\mid\Theta)^{2}+\delta_{\mathrm{\mu_{R}},i}^{2}))^{-1/2}\\ &\times(2\pi(\sigma(m_{i}\mid\Theta)^{2}+\delta_{\mathrm{\mu_{T}},i}^{2}))^{-1/2}\\ &\times\exp\Bigg{(}-\frac{1}{2}\frac{{\mathrm{\mu_{R}},i}^{2}}{\sigma(m_{i}\mid\Theta)^{2}+\delta_{\mathrm{\mu_{R}},i}^{2}}-\frac{1}{2}\frac{{\mathrm{\mu_{T}},i}^{2}}{\sigma(m_{i}\mid\Theta)^{2}+\delta_{\mathrm{\mu_{T}},i}^{2}}\Bigg{)}\end{split}$ (9)
@@
Figure 12 displays the total velocity dispersion as a function of stellar mass for fields A, B, C, and D combined with Gaia XP and Stetson data. The top row shows the fits of the degrees of energy equipartition $\eta$ and $\mu$ for 1G stars, while the second row shows the same for 2G stars. The uncertainties are derived from the 16th and 84th percentiles of the sampled parameter distributions, representing the lower and upper bounds of a 68% credible interval. We present the results separately for each field, as they have different radial coverage, and energy equipartition can vary with radius, exhibiting higher values in the center of the cluster (Bianchini et al. 2018b). The average radius for each field is indicated within the subplots. Each bin contains more than 100 stars, with denser regions containing approximately 400 stars per bin. We note that the bins are displayed for visual comparison only and not for fitting purposes.
@@ 6 Summary and discussion
  • We compared the radial profiles of mean $\mu_{T}$ for 1G and 2G stars by matching Gaia DR3 with the photometric catalog of Lee (2022), which covers an area of $1^{\circ}\times 1^{\circ}$. We did not find significant differences in these profiles. However, our analysis shows that 2G stars have larger values of the ratio of rotational velocity to velocity dispersion indicating differences in the relative strength of rotational to the random motion components.
  • We measured the skewness of the proper motion components and found that 1G stars exhibit higher skewness in $\mu_{\mathrm{T}}$ compared to 2G stars. This result provides further evidence of kinematic differences between the 1G and 2G populations and adds a new constraint for dynamical studies of multiple populations. We briefly discussed various dynamical processes which might affect the skewness of the velocity distribution but further investigations and dedicated numerical simulations are needed to interpret our observational results and understand the dynamical processes responsible for this difference.
@@
  • Deep multi-epoch HST observations allowed us to determine proper motions of white dwarf stars in field A. Since multiple populations could not be identified among white dwarfs, we compared the kinematics of these stars with the overall dynamical profiles of all stars in other evolutionary stages in the cluster. The large uncertainties associated with white dwarfs’ proper motions prevented us from concluding if they are moving as low- or high-mass stars.